A square has a diagnol of 13 inches. What are the dimensions of the square?
Answer by mananth(16946)

diagonal and the sides form a right triangle. Diagonal is the hypotenuse
leg1^2+leg2^2=hyp^2
x^2+x^2=169
2x^2=169
/2
x^2=84.5
x= sqrt(84.5)
side = 9.19 inches